OpenStudy (anonymous):

Find the slope and y-intercept of the line in the given equation. y=9/10 x-1

OpenStudy (anonymous):

y = m x + b m = 9/10 and b = -1

OpenStudy (anonymous):

m=9/10x+-1 9+-1=8 10+-1=9

OpenStudy (anonymous):

8+9=17

OpenStudy (anonymous):

m, the slope, is 9/10 and b, the y intercept, is -1. A plot is attached.

OpenStudy (anonymous):

i'm lost

OpenStudy (anonymous):

m=9/10x+-1 9+-1=8 10+-1=9 8+9=17